The units' digit of a two-digit numeral is 5 more than the tens' digit. The number is 3 times the sum of its digits. Find the numeral.
Answer by Nate(3500)

Number: 10x + y
The units' digit of a two-digit numeral is 5 more than the tens' digit.
y = 5 + x
The number is 3 times the sum of its digits.
10x + y = 3(x + y)
10x + y = 3x + 3y
7x - 2y = 0
Substitute:
7x - 2y = 0
7x - 2(5 + x) = 0
7x - 10 - 2x = 0
5x = 10
x = 2
Substitute:
y = 5 + x
y = 5 + 2
y = 7
Number: 10x + y = 10*2 + 7 = 27